Solve for x.
1/x + 1/x+16 = 1/15
The solutions are x = ________.

Answer by edjones(8007)

1/x + 1/(x+16) = 1/15
15(x+16)+15x=x(x+16) multiply each side by 15x(x+16) to eliminate fractions.
15x+240+15x=x^2+16x
x^2+16x-15x-15x-240=0
x^2-14x-240=0
(x+10)(x-24)=0
x=-10, x=24
.
Ed